OpenStudy (chunkymonkay):

I need help with this practice question. ~Step By Step~ 2x-4y=6

OpenStudy (chunkymonkay):

@Data_LG2

OpenStudy (chunkymonkay):

Determine if the ordered pair is a solution to the equation 2x-4y=6 ordered pair~(-3, -3)

OpenStudy (chunkymonkay):

@StudyGurl14

OpenStudy (bibby):

plug in an x and solve for y

OpenStudy (chunkymonkay):

okay so 2(-3) -4(-3)=6

OpenStudy (chunkymonkay):

-6+12=6 im hoping that im doing this correctly

OpenStudy (bibby):

and then you get 6=6, is that good enough? I'm not sure myself

OpenStudy (chunkymonkay):

yes i think that is the correct way.

OpenStudy (bibby):

I figured you could plug in x=-3 2(-3)-4y=6 -6-4y=6 add 6 to both sides \(-4y=12\) \(y=-3 \checkmark \)

OpenStudy (chunkymonkay):

yes i am 100% sure that is correct thanks a bunch ;)